# Variational Tail Bounds for Norms of Random Vectors and Matrices

### v3

@@ 6 Further Abstraction via Coupling
Let $F\colon\mathbb{R}^{d}\to\mathbb{R}_{\geq 0}$ be a nonnegative function that has a finite moment with respect to the random variable $X\in\mathbb{R}^{d}$. Consider a random variable $Y\geq 0$ that may depend on $X$ to form a coupling $(X,Y)$. With $\mathrm{ess\,supp}(X)$ denoting the essential support of $X$, define
$\nu\left(\operatorname{\mathbb{P}}\left(Y\,\mid X\right)\right)=\nu_{F}\left(\operatorname{\mathbb{P}}\left(Y\,\mid X\right)\right)$ $\stackrel{{\scriptstyle\textup{\tiny def}}}{{=}}\sup_{x\in\mathrm{ess\,supp}(X)}\frac{1}{\operatorname{\mathbb{P}}\left(Y\geq F(X)\,\mid X=x\right)}$
@@
$\left(\operatorname{\mathbb{E}}\left(F^{p}(X)\right)\right)^{1/p}$ $=\inf_{\operatorname{\mathbb{P}}\left(Y\,\mid X\right)}\left(\operatorname{\mathbb{E}}\left(Y^{p}\right)\right)^{1/p}\nu^{1/p}\left(\operatorname{\mathbb{P}}\left(Y\,\mid X\right)\right)\,.$ (24)
@@
$F\left(X\right)$ $\leq\inf_{\begin{subarray}{c}\operatorname{\mathbb{P}}\left(Y\mid X\right)\\ b\geq 0\\ p\geq 1\end{subarray}}b+e^{t/p}\left(\operatorname{\mathbb{E}}\left((Y-b)_{+}^{p}\right)\right)^{1/p}\nu^{1/p}\left(\operatorname{\mathbb{P}}\left(Y\,\mid X\right)\right)\,.$ (25)
@@
Then, (24) follows by setting $b=0$ and taking the infimum with respect to $\operatorname{\mathbb{P}}(Y\,\mid X)$ and observing that for $Y=F(X)$ we have $\nu\left(\operatorname{\mathbb{P}}(Y\,\mid X)\right)=1$. The tail bound (25) follows from (26), Markov’s inequality, and the inequality $F(X)\leq\left(F(X)-b\right)_{+}+b$. ∎
@@
We emphasize that, while the choice of $Y=F(X)$ attains the infimum in (24), we are usually interested in situations where a nearly optimal bound can be achieved by restricted choices of the coupling $(X,Y)$. An obvious practical requirement for suitable couplings is that $\operatorname{\mathbb{E}}\left(Y^{p}\right)$ and $\nu\left(\operatorname{\mathbb{P}}\left(Y\,\mid X\right)\right)$ must be easy-to-approximate. For example, Lemma˜1, as a special case of Lemma˜2 with $F\left(X\right)={\left\lVert X\right\rVert}$, only considers $Y=\left|{\langle U,X\rangle}\right|$ where $U\sim\operatorname{\mathbb{P}}_{0}$ is independent of $X$. It is also worth mentioning that Lemma˜2 uses a slightly refined definition of $\nu\left(\operatorname{\mathbb{P}}_{0}\right)$ of Lemma˜1 in cases where $X/{\left\lVert X\right\rVert}$ is (almost surely) supported on a proper subset of $\partial\mathcal{B}$.
Below we provide a simple example to illustrate the application of Lemma˜2.
@@
Given $\sigma_{1}\geq\dotsc\geq\sigma_{d}>0$, for $i=1,\dotsc,d$ let $X_{i}\sim\mathrm{Normal}(0,\sigma_{i}^{2})$ be independent centered Gaussian random variables. For $p\geq 1$, our goal is to bound $\left(\operatorname{\mathbb{E}}\left({\left\lVert X\right\rVert}_{\infty}^{p}\right)\right)^{1/p}$ where $X=\begin{bmatrix}X_{i}\end{bmatrix}_{i=1}^{d}\in\mathbb{R}^{d}$. Take $J$ to be uniformly distributed over $[d]$ independent of everything else. Choosing $F(X)=\left({\left\lVert X\right\rVert}_{\infty}-\mu\right)_{+}$ and $Y=\left(|X_{J}|-\mu\right)_{+}$ for some parameter $\mu\in\mathbb{R}_{\geq 0}$, we have $\nu(\operatorname{\mathbb{P}}\left(Y\mid\,X\right))\leq d$ and it follows from Lemma˜2 that
$\operatorname{\mathbb{E}}\left(\left({\left\lVert X\right\rVert}_{\infty}-\mu\right)_{+}^{p}\right)$ $\leq d\operatorname{\mathbb{E}}\left(Y^{p}\right)$
$=\sum_{j=1}^{d}\operatorname{\mathbb{E}}\left(\left(|X_{j}|-\mu\right)_{+}^{p}\right)\,.$
@@
$\operatorname{\mathbb{E}}\left(\left(|X_{j}|-\mu\right)_{+}^{p}\right)$ $=\int_{0}^{\infty}\operatorname{\mathbb{P}}\left(|X_{j}|\geq s+\mu\right)ps^{p-1}\mathrm{d}s$
$\leq\int_{0}^{\infty}\inf_{\lambda\geq 0}2e^{\lambda^{2}\sigma_{i}^{2}/2-\lambda(s+\mu)}ps^{p-1}\mathrm{d}s$
$=\int_{0}^{\infty}2e^{-(s+\mu)^{2}/(2\sigma_{i}^{2})}ps^{p-1}\mathrm{d}s$
$\leq 2\sigma_{i}^{p}e^{-\mu^{2}/(2\sigma_{i}^{2})}\left(\frac{p}{2}\right)!\,.$
Therefore, we have
$\left(\operatorname{\mathbb{E}}\left({\left\lVert X\right\rVert}_{\infty}^{p}\right)\right)^{1/p}$ $\leq\mu+\left(\operatorname{\mathbb{E}}\left(\left({\left\lVert X\right\rVert}_{\infty}-\mu\right)_{+}^{p}\right)\right)^{1/p}$
$\leq\mu+\sqrt{2p}\left(\sum_{i=1}^{d}\sigma_{i}^{p}e^{-\mu^{2}/(2\sigma_{i}^{2})}\right)^{1/p}\,.$
Since the optimal choice of $\mu\geq 0$ cannot be expressed explicitly as a function of $p$ and the $\sigma_{i}$s, we resort to approximation. Let
@@
and choose $\mu=c\sigma_{*}$ for some coefficient $c>0$ to obtain
$\left(\operatorname{\mathbb{E}}\left({\left\lVert X\right\rVert}_{\infty}^{p}\right)\right)^{1/p}$ $\leq c\sigma_{*}+\sqrt{2p}\sigma_{1}\left(\sum_{i=1}^{d}\left(i+1\right)^{-c^{2}/2}\right)^{1/p}\,.$
@@
$\left(\operatorname{\mathbb{E}}\left({\left\lVert X\right\rVert}_{\infty}^{p}\right)\right)^{1/p}$ $<2\sigma_{*}+\sqrt{2p}\sigma_{1}\,$,
@@
which, as expected, coincides with the bound obtained from the tight estimate of $\operatorname{\mathbb{E}}\left({\left\lVert X\right\rVert}_{\infty}\right)$ (see, e.g., [vHan17, Section 2.1]) and the Gaussian concentration inequality.
Appendix A A Tail bound for Euclidean norm of sub-exponential random vectors
We use Lemma˜1 to prove the following tail bound for the Euclidean norm of sub-exponential random vectors which is, up to the constant factors, equivalent to the tail bound due to [Zhi24, Proposition 3 ]. As pointed out in [Zhi24], this bound applies to log-concave measures, including uniform measures on convex bodies, and as such it improves on a similar result due to [Ale95]. Also, it can be viewed as a dimension-free specialization of a general result of [LN20] to the case of the Euclidean norm of sub-exponential random vectors.
Proposition 1.
Let $X\in\mathbb{R}^{d}$ be a centered sub-exponential random vector with covariance $\Sigma$, such that for every $u\in\mathbb{R}^{d}$ and $p\geq 1$ we have
$\left(\operatorname{\mathbb{E}}\left(\left|{\langle u,X\rangle}\right|^{p}\right)\right)^{1/p}$ $\leq\eta p{\left\lVert u\right\rVert}_{\Sigma}\,.$
Then for any $t\geq 1$, with probability at least $1-e^{-t}$ we have

### v4

@@ 5 Further Abstraction via Coupling
Let $F\colon\mathbb{R}^{d}\to\mathbb{R}$ be a function that has a finite moment of order $p$ with respect to the random variable $X\in\mathbb{R}^{d}$. Consider a coupling $(X,Y)$ where the random variable $Y\in\mathbb{R}$, which may depend on $X$, is such that $(Y)_{+}$ has a finite moment. With $\mathrm{ess\,supp}(X)$ denoting the essential support of $X$, define
$\nu\left(\operatorname{\mathbb{P}}\left(Y\,\mid X\right)\right)=\nu_{F}\left(\operatorname{\mathbb{P}}\left(Y\,\mid X\right)\right)$ $\stackrel{{\scriptstyle\textup{\tiny def}}}{{=}}\sup_{x\in\mathrm{ess\,supp}(X)}\frac{1}{\operatorname{\mathbb{P}}\left(Y\geq F(X)\,\mid X=x\right)}$ (33)
@@
$\left(\operatorname{\mathbb{E}}\left(\left(F(X)\right)_{+}^{p}\right)\right)^{1/p}$ $=\inf_{\operatorname{\mathbb{P}}\left(Y\,\mid X\right)}\left(\operatorname{\mathbb{E}}\left((Y)^{p}_{+}\right)\right)^{1/p}\nu^{1/p}\left(\operatorname{\mathbb{P}}\left(Y\,\mid X\right)\right)\,.$ (34)
@@
$F\left(X\right)$ $\leq\inf_{\begin{subarray}{c}\operatorname{\mathbb{P}}\left(Y\mid X\right)\\ b\in\mathbb{R}\\ p\geq 1\end{subarray}}b+e^{t/p}\left(\operatorname{\mathbb{E}}\left((Y-b)_{+}^{p}\right)\right)^{1/p}\nu^{1/p}\left(\operatorname{\mathbb{P}}\left(Y\,\mid X\right)\right)\,.$ (35)
More precisely, for any $s\in\mathbb{R}$ we have the following inequality
$\operatorname{\mathbb{P}}\left(F(X)\geq s\right)$ $\leq\nu\left(\operatorname{\mathbb{P}}\left(Y\,\mid\,X\right)\right)\operatorname{\mathbb{P}}\left(Y\geq s\right)\,$, (36)
using which the quantiles of $F(X)$ can be bounded in terms of the quantiles of $Y$.
@@
Then, (34) follows by setting $b=0$ and taking the infimum with respect to $\operatorname{\mathbb{P}}(Y\,\mid X)$ and observing that for $Y=F(X)$ we have $\nu\left(\operatorname{\mathbb{P}}(Y\,\mid X)\right)=1$. The tail bound (35) follows from (37), Markov’s inequality, and the inequality $F(X)\leq\left(F(X)-b\right)_{+}+b$.
To prove (36), observe that
$\operatorname{\mathbb{P}}\left(Y\geq s\right)$ $=\operatorname{\mathbb{E}}\left(\mathbb{1}\left(Y\geq s\right)\right)$
$\geq\operatorname{\mathbb{E}}\operatorname{\mathbb{E}}\left(\mathbb{1}\left(Y\geq F(X)\right)\mathbb{1}\left(F(X)\geq s\right)\,\mid\,X\right)$
$=\operatorname{\mathbb{E}}\left(\operatorname{\mathbb{P}}\left(Y\geq F(X)\,\mid\,X\right)\mathbb{1}\left(F(X)\geq s\right)\right)$
$\geq\operatorname{\mathbb{P}}\left(F(X)\geq s\right)/\nu\left(\operatorname{\mathbb{P}}\left(Y\,\mid\,X\right)\right)\,$,
multiplying both sides by $\nu\left(\operatorname{\mathbb{P}}\left(Y\,\mid\,X\right)\right)$ yields the claimed inequality. ∎
@@
While the choice of $Y=F(X)$ attains the infimum in (34), interesting situations occur when a nearly optimal bound can be achieved by restricted choices of the coupling $(X,Y)$. An obvious practical requirement for suitable couplings is that $\operatorname{\mathbb{E}}\left(Y^{p}\right)$ and $\nu\left(\operatorname{\mathbb{P}}\left(Y\,\mid X\right)\right)$ must be easy-to-approximate. For example, Lemma˜1, as a special case of Lemma˜2 with $F\left(X\right)={\left\lVert X\right\rVert}$, only considers $Y=\left|{\langle U,X\rangle}\right|$ where $U\sim\operatorname{\mathbb{P}}_{0}$ is independent of $X$. A more refined choice is a Gibbs measure with a suitable potential function as the conditional law of $U$ given $X$. It is also worth mentioning that Lemma˜2 uses a slightly refined definition of $\nu\left(\operatorname{\mathbb{P}}_{0}\right)$ of Lemma˜1 in cases where $X/{\left\lVert X\right\rVert}$ is (almost surely) supported on a proper subset of $\partial\mathcal{B}$.
We can also use a similar argument to bound the expectation of supremum over functions of $X$ that belong to a prescribed class of functions. The following proposition provides such a bound.
Proposition 1.
Given an index set $\Theta$ let $\left(F_{\theta}\colon{}\mathbb{R}^{d}\to\mathbb{R}\right)_{\theta\in\Theta}$ be a set of functions that each have a finite moment with respect to the random variable $X\in\mathbb{R}^{d}$. With $(X,Y)\in\mathbb{R}^{d}\times\mathbb{R}$ being a coupling induced by the conditional distribution $\operatorname{\mathbb{P}}\left(Y\,\mid\,X\right)$ such that $(Y)_{+}$ has a finite moment, and recalling the definition of $\nu_{F}(\cdot)$ in (33), we have
$\operatorname{\mathbb{E}}\left(\sup_{\theta\in\Theta}F_{\theta}(X)\right)$ $\leq\inf_{p\geq 1,\,\operatorname{\mathbb{P}}\left(Y\,\mid\,X\right)}\left(\operatorname{\mathbb{E}}\left((Y)_{+}^{p}\right)\right)^{1/p}\sup_{\theta\in\Theta}\nu_{F_{\theta}}^{1/p}(\operatorname{\mathbb{P}}\left(Y\,\mid\,X\right))\,.$ (38)
Proof.
Following the same argument as in the proof of Lemma˜2, for all $\theta\in\Theta$ we have
$\left(F_{\theta}(X)\right)_{+}^{p}$ $\leq\operatorname{\mathbb{E}}\left(\left(Y\right)_{+}^{p}\,\mid\,X\right)\nu_{F_{\theta}}\left(\operatorname{\mathbb{P}}\left(Y\,\mid\,X\right)\right)\,.$
Taking the $p$-th root on both sides and then maximizing with respect to $\theta\in\Theta$ we get
$\sup_{\theta\in\Theta}\left(F_{\theta}(X)\right)_{+}$ $\leq\left(\operatorname{\mathbb{E}}\left(\left(Y\right)_{+}^{p}\,\mid\,X\right)\right)^{1/p}\sup_{\theta\in\Theta}\nu^{1/p}_{F_{\theta}}\left(\operatorname{\mathbb{P}}\left(Y\,\mid\,X\right)\right)\,.$
Then, (38) follows by taking the expectation with respect to $X$, applying Jensen’s inequality for the concave function $z\mapsto z^{1/p}$, and using the fact that $\sup_{\theta\in\Theta}(F_{\theta}(x))_{+}\geq\sup_{\theta\in\Theta}F_{\theta}(x)$ for all $x\in\mathbb{R}^{d}$. ∎
[…]
$D_{\alpha}\left(\mu,\mu_{\mathrm{ref}}\right)$ $=\frac{1}{\alpha-1}\log\operatorname{\mathbb{E}}_{\mu_{\mathrm{ref}}}\left(\left(\frac{\mathrm{d}\mu}{\mathrm{d}\mu_{\mathrm{ref}}}\right)^{\alpha}\right)\,.$
Corollary 2.
Let $X\sim\mu_{X}$ be a random variable in $\mathbb{R}^{d}$, $\mathcal{M}$ being a set of probability measures on some sample space $\mathcal{Z}$, and $f\colon{}\mathbb{R}^{d}\times\mathcal{Z}\to\mathbb{\mathbb{}}R$ be a function that for each $\mu\in\mathcal{M}$ has a finite $L^{p}(\mu_{X}\times\mu)$ norm for some $p\geq 1$. Denote the median [^7] We chose the median instead of any other fixed quantile for simplicity and to avoid introducing extra parameters. of $f(x,Z)$ with respect to $Z\sim\mu$ by
$F_{\mu}(x)=\operatorname*{med}_{Z\sim\mu}\left(f(x,Z)\right)\,.$
Then, for any reference measure $\mu_{\mathrm{ref}}\in\mathcal{M}$ and $\alpha>1$ we have
$\operatorname{\mathbb{E}}_{\mu_{X}}\left(\sup_{\mu\in\mathcal{M}\colon{}\mu\ll\mu_{\mathrm{ref}}}\hskip-8.61108ptF_{\mu}(X)\right)$ $\leq\inf_{p\geq 1}\left(\left(\operatorname{\mathbb{E}}_{\mu_{X}\times\mu_{\mathrm{ref}}}\left(\left(f(X,Z)\right)_{+}^{p}\right)\right)^{1/p}\sup_{\mu\in\mathcal{M}\colon{}\mu\ll\mu_{\mathrm{ref}}}\hskip-8.61108pte^{\frac{1}{p}D_{\alpha}(\mu,\mu_{\mathrm{ref}})+\frac{\alpha\log 2}{(\alpha-1)p}}\right)\,.$
Proof.
For any $\mu_{\mathrm{ref}}$-measurable event $\mathcal{E}$, it follows from the definition of the Rényi divergence for $\alpha>1$ and Hölder’s inequality that
$\mu(\mathcal{E})$ $=\operatorname{\mathbb{E}}_{\mu_{\mathrm{ref}}}\left(\frac{\mathrm{d}\mu}{\mathrm{d}\mu_{\mathrm{ref}}}\cdot\mathbb{1}_{\mathcal{E}}\right)$
$\leq\left(\operatorname{\mathbb{E}}_{\mu_{\mathrm{ref}}}\left(\left(\frac{\mathrm{d}\mu}{\mathrm{d}\mu_{\mathrm{ref}}}\right)^{\alpha}\right)\right)^{1/\alpha}\left(\operatorname{\mathbb{E}}_{\mu_{\mathrm{ref}}}\left(\mathbb{1}_{\mathcal{E}}\right)\right)^{(\alpha-1)/\alpha}$
$=e^{(\alpha-1)D_{\alpha}(\mu,\mu_{\mathrm{ref}})/\alpha}\left(\mu_{\mathrm{ref}}(\mathcal{E})\right)^{(\alpha-1)/\alpha}\,.$
In particular, for $\mathcal{E}={\mathcal{E}}_{x}=\{f(x,Z)\geq F_{\mu}(x)\}$, we get
$\operatorname{\mathbb{P}}_{Z\sim\mu_{\mathrm{ref}}}\left(f(x,Z)\geq F_{\mu}(x)\right)$ $\geq\left(\operatorname{\mathbb{P}}_{Z\sim\mu}\left(f(x,Z)\geq F_{\mu}(x)\right)\right)^{\alpha/(\alpha-1)}e^{-D_{\alpha}(\mu,\mu_{\mathrm{ref}})}$
$\geq 2^{-\alpha/(\alpha-1)}e^{-D_{\alpha}(\mu,\mu_{\mathrm{ref}})}\,$,
where the second line follows from the definition of $F_{\mu}(\cdot)$. The claim follows from Proposition˜1 with $Y=f(X,Z)$ for $Z\sim\mu_{\mathrm{ref}}$, treating $\mu$ as the index for $F_{\mu}(\cdot)$, and using the inequality above which guarantees $\nu_{F_{\mu}}\left(\operatorname{\mathbb{P}}\left(Y\,\mid\,X\right)\right)\leq 2^{\alpha/(\alpha-1)}e^{D_{\alpha}(\mu,\mu_{\mathrm{ref}})}$. ∎
Below we provide a simple application of Proposition˜1.
@@
Given $\sigma_{1}\geq\dotsc\geq\sigma_{d}>0$, for $i=1,\dotsc,d$ let $X_{i}\sim\mathrm{Normal}(0,\sigma_{i}^{2})$ be independent centered Gaussian random variables. Our goal is to bound $\operatorname{\mathbb{E}}\left({\left\lVert X\right\rVert}_{\infty}\right)$ where $X=\begin{bmatrix}X_{i}\end{bmatrix}_{i=1}^{d}\in\mathbb{R}^{d}$. Let $\Theta=[d]$ and for $\theta\in\Theta$ define $F_{\theta}\colon{}\mathbb{R}^{d}\to\mathbb{R}$ as $F_{\theta}(x)=|x_{\theta}|-b$ for some parameter $b\geq 0$. Furthermore, take $J$ to be uniformly distributed over $[d]$ independent of everything else and let $Y=|X_{J}|-b$. We have $\nu_{F_{\theta}}(\operatorname{\mathbb{P}}\left(Y\mid\,X\right))\leq d$ and it follows from Proposition˜1 that
$\operatorname{\mathbb{E}}\left(\max_{i\in[d]}|X_{i}|-b\right)$ $\leq d^{1/p}\left(\operatorname{\mathbb{E}}\left((Y)_{+}^{p}\right)\right)^{1/p}$
$=\left(\sum_{j=1}^{d}\operatorname{\mathbb{E}}\left(\left(|X_{j}|-b\right)_{+}^{p}\right)\right)^{1/p}\,.$
@@
$\operatorname{\mathbb{E}}\left(\left(|X_{j}|-b\right)_{+}^{p}\right)$ $=\int_{0}^{\infty}\operatorname{\mathbb{P}}\left(|X_{j}|\geq s+b\right)ps^{p-1}\mathrm{d}s$
$\leq\int_{0}^{\infty}\inf_{\lambda\geq 0}2e^{\lambda^{2}\sigma_{i}^{2}/2-\lambda(s+b)}ps^{p-1}\mathrm{d}s$
$=\int_{0}^{\infty}2e^{-(s+b)^{2}/(2\sigma_{i}^{2})}ps^{p-1}\mathrm{d}s$
$\leq 2\sigma_{i}^{p}e^{-b^{2}/(2\sigma_{i}^{2})}\left(\frac{p}{2}\right)!\,.$
Combining the derived inequalities and optimizing with respect to $p\geq 1$ we obtain
$\operatorname{\mathbb{E}}\left({\left\lVert X\right\rVert}_{\infty}\right)$ $\leq b+\inf_{p\geq 1}\sqrt{2p}\left(\sum_{i=1}^{d}\sigma_{i}^{p}e^{-b^{2}/(2\sigma_{i}^{2})}\right)^{1/p}\,.$
Since the optimal choice of $b\geq 0$ and $p\geq 1$ cannot be expressed explicitly as a function of the $\sigma_{i}$s, we resort to approximation. Let
@@
and choose $b=c\sigma_{*}$ for some coefficient $c>0$ to obtain
$\operatorname{\mathbb{E}}\left({\left\lVert X\right\rVert}_{\infty}\right)$ $\leq c\sigma_{*}+\inf_{p\geq 1}\sqrt{2p}\sigma_{1}\left(\sum_{i=1}^{d}\left(i+1\right)^{-c^{2}/2}\right)^{1/p}\,.$
@@
$\operatorname{\mathbb{E}}\left({\left\lVert X\right\rVert}_{\infty}\right)$ $<2\sigma_{*}+\sqrt{2}\sigma_{1}\,$,
@@
which, in view of the fact that $\sigma_{1}\leq\sigma_{*}/\sqrt{\log 2}$, coincides with the tight estimate of $\operatorname{\mathbb{E}}\left({\left\lVert X\right\rVert}_{\infty}\right)$ (see, e.g., [52, Section 2.1]). Applying the Gaussian concentration inequality we also obtain the moment and tail bounds for $\left\lVert X\right\rVert$$_{\infty}$.
Appendix A Recovering Example˜2 and (13) via Theorem˜2
We can recover the bound (9) in Example˜1 from Theorem˜2 up to a factor of order $\sqrt{\log c_{k}}$ that can be ignored if $c_{k}$ is not large relative to $\pi_{k}^{-1}$. The parameters of the problem in this example are $\eta_{1}=\sqrt{2}$, $\eta_{2}=0$, and $\Sigma=I$. By taking $\kappa_{0}\to\infty$, we have $f_{0}(x)=\rho_{0}{\left\lVert x\right\rVert}$ which corresponds to the choice
we have
$\geq 2\pi_{k}\,$,
using which we obtain the bound
$\overline{\nu}_{\Sigma}(\infty,2c_{k})$ $\leq\frac{c_{k}}{\pi_{k}}\,.$
Therefore, the bound (15) provided by Theorem˜2 implies
${\left\lVert X\right\rVert}$ $\leq 2\max_{i\in[N]}{\left\lVert u_{i}\right\rVert}^{2}_{2}\inf_{p\geq 1,\,k\in[N]}e^{t/(2p)}\sqrt{2p}c_{k}\left(\frac{c_{k}}{\pi_{k}}\right)^{1/(2p)}$
Let $k_{\star}\in[N]$ be the minimizer of $\min_{k\in[N]}c_{k}\max\left\{1,\sqrt{t-\log\pi_{k}}\right\}$ in (9). If we have $\log c_{k_{\star}}\lesssim t-\log\pi_{k_{\star}}$, our derivations confirm that, up to constant factors, Theorem˜2 recovers the bound (9) of Example˜1.
Next we show that the tail bound (13), which applies when $X$ has sub-Gaussian marginals, can be recovered from Theorem˜2. We have $\eta_{1}=\eta$ and again $\eta_{2}=0$. Also, recall that in this example $\mathcal{B}$ and $\mathcal{B}_{*}$ coincide with the unit $\ell_{2}$-ball. By taking $\rho_{0}\to+\infty$ we have $f_{0}(x)=\kappa_{0}{\left\lVert x\right\rVert}_{\Sigma^{-1/2}}^{2}/2$, and thus
$U=\nabla f_{0}(\Sigma^{1/2}G)$ $=\kappa_{0}G\,.$